I need major help with these two questions ASAP!
1.) The sum of two numbers is 78. Three times the lesser number is six less than the greater number. What are the numbers?
2.) The sum of three consecutive odd numbers is 273. What are the numbers?

Thanks whoever solves this problem it helps so much!

Answer by stanbon(57282)

1.) The sum of two numbers is 78. Three times the lesser number is six less than the greater number. What are the numbers?
Let the numbers be x and y.
EQUATIONS:
x+y = 78
3x=y-6
---------
Rewrite the equations
y = -x+78
y = 3x+6
------------
Substitute and solve for "x":
-x+78 = 3x+6
4x = 72
x = 18
--------
Substitute to solve for y:
18+y = 78
y = 60
============
2.) The sum of three consecutive odd numbers is 273. What are the numbers?
Let the integers be x+1, x+3, x+5
EQUATION:
x+1 + x+3 + x+5 = 273
3x +9 = 273
3x = 264
x = 88
---------
x+1 = 89
x+3 = 91
x+5 = 93
-----------
Cheers,
Stan H.